# Mosaic IT: Enhancing Instruction Tuning with Data Mosaics

### v1

@@ A.1 Distribution description
where $\mu=0$ and $\sigma=0$ by default in our setting. We will resample with this distribution if the sampled value $x_{sample}$ is greater then $k_{max}$.
@@
where $\mu=0$ and $s=2$ by default in our setting. We will resample with this distribution if the sampled value $x_{sample}$ is greater then $k_{max}$.
@@
where $m=1$ and $\alpha=1$ by default in our setting. We will resample with this distribution if the sampled value $x_{sample}-1$ is greater then $k_{max}$.
@@ A.2 Distribution visualization
The detailed data counts for different distribution are provided in Figure 3.
@@ B.1 Implementation Details
The experiments are conducted on three base pre-trained models: Llama2-7B, Llama2-13B [46], and Mistral-7B [19]. We utilize the prompt and code base from Vicuna [7] and flash attention [10] while the overall training arguments are aligned with the common training configuration. The Adam optimizer [21] is utilized with the batch size of $128$ and with the max token length of $2048$. When training the baseline models on Llama2-7B and Llama2-13B, the maximum learning rate is set to $2\times 10^{-5}$ with the warmup rate as $0.03$ for $3$ epochs. When training the baseline models on Mistral-7B, the maximum learning rate is set to $1\times 10^{-5}$ with the warmup rate as $0.1$ for $3$ epochs. When training with Mosaic-IT, we run the mosaic process $4$ times for each experiment to simulate $4$ epochs of training. Then these augmented data are mixed together used for training $1$ epoch while all other configurations are kept the same as baselines.
@@ B.2 Training Dataset
The Alpaca dataset [44] comprises $52,000$ instruction-following samples and is constructed utilizing the self-instruct paradigm [49]. This dataset was produced by employing OpenAI’s text-davinci-003 model. Characterized as a classical dataset with moderate quality attributes, the Alpaca dataset serves as an initial platform to validate our methodology. To further substantiate our approach using a dataset of inherently high quality, we also applied our method to the Alpaca-GPT4 dataset [40], which features responses generated by GPT4. The WizardLM dataset [54] is also utilized in our method, which contains $70,000$ samples created by the evolution algorithm proposed by them. With ChatGPT-3.5 utilized, the data quality on WizardLM is largely guaranteed.
@@ B.3 Evaluation Metrics
Open LLM Leaderboard, employing the evaluation framework from Eval Harness [16], offers a detailed and systematic approach to assessing the capabilities of generative language models through a set of diverse evaluation tasks. This methodology zeroes in on four pivotal benchmarks: ARC [9], HellaSwag [60], MMLU [18], and TruthfulQA [30]. These benchmarks collectively provide a comprehensive evaluation of the models’ reasoning abilities, their grasp of common-sense knowledge, and their accuracy in presenting factual information. Consequently, the leaderboard presents valuable insights.
@@
Alpaca-Eval Leaderboard, leveraging the AlpacaFarm evaluation dataset, presents a dependable and efficient automated evaluation tool for LLMs [29, 15]. This tool benchmarks the responses generated by LLMs against those from Davinci003, focusing on the models’ ability to adhere to generic user instructions.
@@
MT-Bench (Multi-turn Benchmark) [66] is a benchmark tool designed for automated evaluating LLMs in multi-turn dialogue settings. It focuses on analyzing conversation flow and the model’s ability to follow instructions with 80 high-quality, multi-turn questions.
@@
IFEval (Instruction-Following Eval) [61] is a straightforward and easy-to-produce evaluation benchmark focusing on a set of “verifiable instructions”. It contains 25 types of verifiable instructions and 541 prompts, with each prompt containing one or multiple verifiable instructions.
Appendix C Predefined Rules
@@
[i] <text> RESPONSE, END [1]. <RESPONSE>response<END>
<i> <<text>> RESPONSE, END OF RESPONSE <1>. <<RESPONSE>>response<<END OF RESPONSE>>
<<i>> |text| OPEN, CLOSE <<1>>. |OPEN|response|CLOSE|
@@
##i <|text|> INITIATE, TERMINATE ##1. <|INITIATE|>response<|TERMINATE|>
##i## #text# START POINT, END POINT ##1##. #START POINT#response#END POINT#
@@
Appendix D Prompt for Evaluation

### v2

@@ A.1 Distribution description
where $\mu=0$ and $\sigma=0$ by default in our setting. We will resample with this distribution if the sampled value $x_{sample}$ is greater than $k_{max}$.
@@
where $\mu=0$ and $s=2$ by default in our setting. We will resample with this distribution if the sampled value $x_{sample}$ is greater than $k_{max}$.
@@
where $m=1$ and $\alpha=1$ by default in our setting. We will resample with this distribution if the sampled value $x_{sample}-1$ is greater than $k_{max}$.
@@ A.2 Distribution visualization
The detailed data counts for different distributions are provided in Figure 5.
@@ B.1 Implementation Details
For the three base pre-trained models, Llama2-7B, Llama2-13B Touvron et al. (2023b), and Mistral-7B Jiang et al. (2023), we utilize the prompt and code base from Vicuna Chiang et al. (2023) and flash attention Dao et al. (2022). The overall training arguments are aligned with the common training configuration. The Adam optimizer Kingma & Ba (2017) is utilized with the batch size of $128$ and with the max token length of $2048$. When training the baseline models Llama2-7B and Llama2-13B, the maximum learning rate is set to $2\times 10^{-5}$ with the warmup rate as $0.03$ for $3$ epochs. When training the baseline models on Mistral-7B, the maximum learning rate is set to $1\times 10^{-5}$ with the warmup rate as $0.1$ for $3$ epochs. For the three models, Llama-3-8B Dubey et al. (2024), Phi-3 Abdin et al. (2024), and Gemma2-2B Team et al. (2024), we utilize the code base from LLaMA-Factory Zheng et al. (2024b). The max token length is set with $4096$ following the modern settings and we train the model for $2$ epochs. Other parameters are kept the same as the above.
When training with Mosaic-IT, we run the mosaic process $n$ times for each experiment to simulate $n$ epochs of training, $n$ represents the number of epochs trained on baseline models, to ensure the alignment of overall data sample counts. Then these augmented data are mixed together and used for training $1$ epoch while all other configurations are kept the same as baselines.
@@ B.2 Training Dataset
The Alpaca dataset Taori et al. (2023) comprises $52,000$ instruction-following samples and is constructed utilizing the self-instruct paradigm Wang et al. (2023b). This dataset was produced by employing OpenAI’s text-davinci-003 model. Characterized as a classical dataset with moderate quality attributes, the Alpaca dataset serves as an initial platform to validate our methodology. To further substantiate our approach using a dataset of inherently high quality, we also applied our method to the Alpaca-GPT4 dataset Peng et al. (2023), which features responses generated by GPT4. The WizardLM dataset Xu et al. (2023) is also utilized in our method, which contains $70,000$ samples created by the evolution algorithm proposed by them. With ChatGPT-3.5 utilized, the data quality on WizardLM is largely guaranteed. The Vicuna 1M dataset Zheng et al. (2024a) is a large-scale dataset containing one million real-world conversations with 25 state-of-the-art LLMs, due to the computation budget, $300$k instances are randomly sampled for our experiments. Magpie dataset Xu et al. (2024b) is a most recent SOTA synthetic dataset with $300$k samples.
@@ B.3 Evaluation Metrics
Pair-wise Comparison by using powerful LLMs like GPT-4 is recently widely accepted and becoming a common practice Touvron et al. (2023b); Chiang et al. (2023); Dettmers et al. (2023); Liu et al. (2023b); Chiang & Lee (2023). The evaluation of responses from LLMs, especially in open-domain contexts where definitive ground truth is hard to establish, continues to be an intricate and evolving research domain. Recent studies, however, have indicated a notable alignment between GPT-4’s performance evaluations and human assessments Zheng et al. (2023); Li et al. (2023c); Sottana et al. (2023), thereby establishing a credible foundation for this evaluative methodology. We adopted test instruction sets from WizardLM Xu et al. (2023), comprising 218 diverse, human-curated instructions for pair-wise comparison. We directly follow the evaluation framework proposed by Chen et al. (2023); Li et al. (2024d), which evaluates responses on a scale spanning from 1 to 10 across multiple dimensions. To further address positional bias, as discussed by Ko et al. (2020); Wang et al. (2023a), the comparison is conducted in two distinct sequences, LLM1’s response first and then LLM2’s response first, ensuring a fair assessment of model performance. Evaluation outcomes are categorized into ’win-tie-loss’ for each instruction. The detailed evaluation prompt is provided in Appendix E.
@@
Open LLM Leaderboard, employing the evaluation framework from Eval Harness Gao et al. (2021), offers a detailed and systematic approach to assessing the capabilities of generative language models through a set of diverse evaluation tasks. This methodology zeroes in on four pivotal benchmarks: ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2021), and TruthfulQA Lin et al. (2022). These benchmarks collectively provide a comprehensive evaluation of the models’ reasoning abilities, their grasp of common-sense knowledge, and their accuracy in presenting factual information. Consequently, the leaderboard presents valuable insights.
@@
Alpaca-Eval Leaderboard, leveraging the AlpacaFarm evaluation dataset, presents a dependable and efficient automated evaluation tool for LLMs Li et al. (2023c); Dubois et al. (2023). This tool benchmarks the responses generated by LLMs against those from Davinci003, focusing on the models’ ability to adhere to generic user instructions.
@@
MT-Bench (Multi-turn Benchmark) Zheng et al. (2023) is a benchmark tool designed for automated evaluating LLMs in multi-turn dialogue settings. It focuses on analyzing conversation flow and the model’s ability to follow instructions with 80 high-quality, multi-turn questions.
@@
IFEval (Instruction-Following Eval) Zeng et al. (2024) is a straightforward and easy-to-produce evaluation benchmark focusing on a set of “verifiable instructions”. It contains 25 types of verifiable instructions and 541 prompts, with each prompt containing one or multiple verifiable instructions.
Appendix C Related Work
Earlier research in instruction tuning primarily centered on constructing expansive, high-quality datasets through intensive curation by human experts, a process both time-consuming and labor-intensive Khashabi et al. (2020); Ye et al. (2021); Wei et al. (2022); Wang et al. (2022); Du et al. (2022). Motivated by the success of Alpaca Taori et al. (2023), recent studies have explored automated approaches for developing instruction-tuning datasets.
Instruction Data Improvement: WizardLM Xu et al. (2023) first proposes an Evol Algorithm to complicate the existing data and reach supreme performance. LaMini-LM Wu et al. (2024) innovatively generates ”Topic-Guided” instructions utilizing Wiki data. Tree-Instruct Zhao et al. (2024) preliminarily explores the relationship between instruction complexity and Alignment and proposes adding nodes to complicate the instruction. UltraChat Ding et al. (2023) establishes broad thematic scopes, systematically generating numerous instructions within each. Reflection-Tuning Li et al. (2023a) sequentially refines both instructions and responses by focusing on specific evaluative criteria. DEITA Liu et al. (2023a) utilizes ChatGPT to diversify and then select the data. Selective Reflection-Tuning Li et al. (2024b) proposes a teacher-student collaborative pipeline to improve and select the data. Instruction Fusion Guo et al. (2024) proposes to utilize ChatGPT4 to merge two distinct instructions for further complexity enhancement. These advancements showcase a shift towards automating the generation and refinement of datasets, reducing reliance on human labor.
Instruction Data Selection: It is widely accepted that ”quality is all you need” Touvron et al. (2023b); Zhou et al. (2023) for instruction tuning. LIMA Zhou et al. (2023) demonstrates that merely 1,000 human-carefully-curated, high-quality training instances can substantially enhance the instruction-following performance. InsTag Lu et al. (2023) employs the proprietary model, ChatGPT, to tag instruction data and select data with complex tags. Alpagasus Chen et al. (2023) utilizes proprietary LLMs chatGPT and Claude2 to directly assess the quality of instruction tuning data. Cherry LLM Li et al. (2024d) proposes the Instruction-Following Difficulty (IFD) scores to assess the difficulty of the instructions, which is a self-guided method in which no extra LLMs are utilized. Motivated by Humpback Li et al. (2023b), Selective Reflection-Tuning Li et al. (2024b) extends the IFD score to a reverse version, focusing on the feasibility of responses. Du et al. (2023) and Bukharin & Zhao (2023) utilize reward models as the base scores for measuring data quality. DEITA Liu et al. (2023a) experiments on several different data selection metrics and builds a dataset with high quality. Superfiltering Li et al. (2024c) reveals the consistency between weak and strong language models in perceiving instruction difficulty, making the filtering process much more efficient. All these works are devoted to distinguishing and selecting good data samples from bad ones for instruction tuning.
Appendix D Predefined Rules
@@
[…]
@@
[…]
##i## ##\##text##\## START POINT, END POINT ##1##. ##\##START POINT##\##response##\##END POINT##\##
@@
Appendix E Prompt for Evaluation
@@
Appendix F Detailed Performance Scores on Llama3, Phi3 and Gemma2
The detailed performance scores on the Open LLM Leaderboard and IFEval, for Llama-3-8B, Phi-3, and Gemma2-2B.
Table 10: The performance comparison on more model families and datasets on all five automatic evaluation metrics. In IF Eval, P and I represent Prompt-level and Instruction-level accuracy.
Model Dataset Method Open LLM Leaderboard $\uparrow$ IF Eval $\uparrow$
Average ARC HellaSwag MMLU TruthfulQA Prompt (S) Inst (S) Prompt (L) Inst (L)
Llama-3-8B Vicuna Baseline 52.51 44.54 70.66 49.68 45.18 19.04 30.70 21.26 33.45
Mosaic-IT 55.62 47.78 73.77 56.11 44.83 29.76 43.17 31.42 45.56
Magpie Baseline 56.15 50.09 71.29 54.40 48.84 29.39 40.76 35.67 47.72
Mosaic-IT 60.13 53.58 76.62 60.82 49.52 38.08 49.64 40.67 52.76
Phi-3 Vicuna Baseline 62.06 58.96 76.48 64.89 47.89 28.47 40.29 30.50 43.17
Mosaic-IT 62.30 58.45 77.66 65.24 47.87 30.13 39.57 32.35 41.85
Magpie Baseline 62.90 59.30 75.07 65.89 51.35 39.56 50.84 44.36 55.25
Mosaic-IT 63.54 60.23 76.30 66.14 51.50 42.33 53.60 50.83 62.35
Gemma2-2B Vicuna Baseline 48.90 43.43 64.20 41.50 46.46 20.51 32.61 23.66 35.61
Mosaic-IT 51.31 46.33 69.32 44.29 45.31 21.44 33.57 24.03 36.93
Magpie Baseline 46.37 39.59 60.71 35.46 49.75 19.78 29.74 21.81 32.49
Mosaic-IT 48.36 39.33 64.10 39.87 50.16 19.78 31.65 22.18 34.77